What is the ending balance of retained earnings for Zippy Company?

Net Income, $40,000

Beginning retained earnings, $700,000

Dividend payout ratio, 30 %
Answer #1

Retained earnings = 1-Dividend payout ratio

Retained earnings = 1-30% = 70%

ending balance of retained earnings for Zippy Company = Beginning retained earnings + New retained earnings

Emding balance of retained earnings = 700000 + 0.7*40000 = $728,000